# XTE J1814-338: A potential hybrid star candidate

### v2

@@ Abstract
So far we have examined the stability of XTE J1814-338 compatible configurations considering that the transition mass is at the $2M_{\odot}$. However, there have been recent measurements that suggest the existence of even more massive compact stars. For instance, Romani et al. Romani-2022 , reported that PSR J0952-0602, one of the fastest rotating pulsars, has a mass of $2.35\pm 0.17M_{\odot}$. For the latter reason, we wish to examine if the hybrid configurations (appearing in Fig. 3), that are compatible with XTE J1814-338, are also stable. In Table 1 we report the properties of the last slow stable configuration for different transition masses (terminal configurations). Beyond these configurations (i.e., for larger central pressure) the eigenvalue of the fundamental mode becomes negative and therefore the derived stellar models are no longer stable. As is evident, the terminal configurations are within the predictions for the XTE J1814-338 properties and therefore the existence of such an object is potentially compatible with the observation of massive compact stars. Notably, the results in Table 1 were derived by considering the DD2 EOS for the description of the hadronic phase. We found that in the case where the low density phase is modeled via the APR EOS, the terminal configurations for $M_{\mathrm{tr}}=2.1,$ $2.15M_{\odot}$ appear before the XTE J1814-338 region is reached.
@@
Summary. In this work we have argued on the explanation of the recently announced XTE J1814-338 properties through the occurence of an extremely strong phase transition in dense matter. In particular, we have constructed hybrid EOSs by considering interestingly large values for the energy density jump. For the description of the hadronic phase we have used the GRDF-DD2 and APR EOSs, while for quark matter we have utilized the widely employed constant speed of sound ansantz.
@@
By analysing the $M-R$ diagrams for different parametrizations of the quark phase, we were able to find hybrid configurations that satisfy the XTE J1814-338 constraints. When considering the GRDF-DD2 model (which is stiffer compared to APR), those stellar models were also stable in terms of the well-known turning point criterion. This was not the case for the APR EOS. However, by performing a complementary analysis, based on the radial oscillations of compact stars, we were able to show that hybrid configurations (compatible with XTE J1814-338) involving the APR model may be stable in a slow conversion scenario. This is of particular importance, as XTE J1814-338, in combination with future observations (that will constrain the stiffness of the nuclear EOS), may not only provide evidence of a strong QCD phase transition but also shed light on the phase conversion dynamics.
@@
Acknowledgements. The authors would like to thank Professor K.D. Kokkotas for his useful insight and comments. The research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 5th Call for HFRI PhD Fellowships (Fellowship Number: 19175).

### v3

@@ IV Results and Discussion
So far we have discussed the stability of XTE J1814-338 compatible configurations considering that the transition mass is at the $2M_{\odot}$. However, there have been recent measurements that suggest the existence of even more massive compact stars. For instance, Romani et al. Romani-2022 , reported that PSR J0952-0602, one of the fastest rotating pulsars, has a mass of $2.35\pm 0.17M_{\odot}$. For the latter reason, we wish to examine if the hybrid configurations (appearing in Fig. 3), which are compatible with XTE J1814-338, are also stable. In Table 1 we report the properties of the last slow stable configuration for different transition masses [marked with x in Fig. 3(b)]. As is evident, the terminal configurations are within the predictions for the XTE J1814-338 properties and therefore the existence of such an object is potentially compatible with the observation of massive compact stars. Notably, the results in Table 1 were derived by considering the DD2 EOS for the description of the hadronic phase. We found that in the case where the low density phase is modeled via the APR EOS, the terminal configurations for $M_{\mathrm{tr}}=2.1,$ $2.15M_{\odot}$ appear before the XTE J1814-338 region ($1\sigma$) is reached. Nonetheless, for $M_{\mathrm{tr}}=2.1M_{\odot}$ there are stable configurations that are compatible with XTE J1814-338 at the $2\sigma$ level (see Fig. 3(a)).
V Summary
@@
In this work we have argued on the explanation of the recently announced XTE J1814-338 properties through the occurrence of an extremely strong phase transition in dense matter. In particular, we have constructed hybrid EOSs by considering interestingly large values for the energy density jump. For the description of the hadronic phase we have used the GRDF-DD2 and APR EOSs, while for quark matter we have utilized the widely employed constant speed of sound ansatz.
@@
By analyzing the $M-R$ diagrams for different parametrizations of the quark phase, we were able to find hybrid configurations that satisfy the XTE J1814-338 constraints. When considering the GRDF-DD2 model (which is stiffer compared to APR), those stellar models were also stable in terms of the well-known turning point criterion. This was not the case for the APR EOS. However, by performing a complementary analysis, based on the radial oscillations of compact stars, we were able to show that hybrid configurations (compatible with XTE J1814-338) involving the APR model may be stable in a slow conversion scenario. This is of particular importance, as XTE J1814-338, in combination with future observations (that will constrain the stiffness of the nuclear EOS), may not only provide evidence of a strong QCD phase transition but also shed light on the phase conversion dynamics.
@@
Acknowledgements.
The authors would like to thank Professor K.D. Kokkotas for his useful insights and comments and Mr Y. Kini for providing the data for the contour regions for XTE J1814-338. The research work was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 5th Call for HFRI PhD Fellowships (Fellowship No. 19175).